OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 - 8x - 42.

OpenStudy (anonymous):

first factor out a 2, then.......

OpenStudy (anonymous):

Consider a general quadratic equation written as follows: ax^2 +bx + c where a, b and c are constants. one quick way of factorizing this is to split the bx term using the following criteria: find two numbers d and h such that d+h = b and d*h = a*c => ax^2 + bx +c = ax^2 + (d+h)x + c (i.e replace b with (d+h)) ax^2 + bx +c = ax^2 + (d+h)x + c= ax^2 + dx + hx + c factorize the above equation and group the factors. For your example: If you first factor 2 as suggested by lag: then rewrite the equation as follows: 2x^2-8x - 42 = 2*(x^2 -4x -21) Factorize the equation in the brackets by comparing with ax^2 + bx +c: x^2 -4x -21 => b = -4, a = 1, c = -21 choose d= -7 and h =3 i.e. d+h = -7+3 = -4 = b and d*h = -7*3 = -21 = C so re-write x^2 -7x -21 = x^2 -7x + 3x -21 factorize x^2 -7x + 3x -21 = x(x-7) + 3(x-7) group: the factors (x+3)(x-7) hence x^2-7x-21 = (x+3)(x-7) since you had factored 2 originally the final answer is 2*(x+3)(x-7)